OpenStudy (anonymous):

2. Factor completely: 3x2 – 21x + 30 (Points : 1) (x – 2)(x – 5) (3x – 2)(x – 15) 2(x – 3)(x – 10) 3(x – 2)(x – 5) Question 3. 3. Factor completely: 3x3 − 48x (Points : 1) x(x − 4)(x − 4) 3x(x − 16) 3x(x + 4)(x − 4) 3(x + 4)(x − 4)

OpenStudy (mrnood):

start by looking at the answers the last terms multiplied toghether must equal the last term in the equation so in the first 1 option a: -2 x -5 = 10 so it cannot be the answer the first terms multipled together must equal the fist term in th eequation so again in the first option x *x = x^2 but in the equation you have 3x^2 Any that meet the above - just multiply them out to see if they come to the original